# DONOD: Robust and Generalizable Instruction Fine-Tuning for LLMs via Model-Intrinsic Dataset Pruning

### v1

@@ 4.1 Experiment Setup
Following Ma et al. (2025), we constructed our evaluation benchmark based on AGIEval from Zhong et al. (2024) and IFEval from Zhou et al. (2023). This benchmark is designed to be comprehensive and domain-orthogonal, measuring abilities in logical reasoning, mathematics, reading comprehension, and instruction following. By incorporating diverse datasets, the benchmark closely mirrors real-world scenarios of ad-hoc SFT, where the goal is to enhance specific abilities of the model rather than simply improving performance on a limited and unrepresentative benchmark. The details of the benchmark are shown in Table A.1.
For training datasets, we experimented with DONOD on two different datasets to demonstrate its dataset-agnostic nature and applicability across domains, regardless of the availability of a validation set. For mathematical problem-solving, we chose the SAT Math Chain-of-Thought (COT) dataset from Davidson (2023), which is constructed from the real-world Scholastic Assessment Test (SAT) question test-bank. This dataset provides no validation set, but its distribution closely aligns with the SAT-math test set. We show that DONOD generalizes well not only to closely related distributions but also to other math domains and unrelated tasks. Conversely, we also selected the LogiQA-Train dataset from Liu et al. (2020), which includes both a training set and a validation set. Unlike the SAT Math dataset, LogiQA is not COT labeled, providing only the correct option letter as the answer for each multiple-choice question. The details of the training datasets are shown in Table  5.
Models and Experiment Settings
DONOD is evaluated on three models: Llama-3.2-3B-Instruct Meta AI (2024), Llama-3.1-8B-Instruct Grattafiori et al. (2024), and Llama-2-13b-chat Touvron et al. (2023). These models were chosen to demonstrate the generalization of DONOD across different architectures and fine-tuning paradigms: Llama-3.2-3B-Instruct is a smaller model optimized for instruction following, while Llama-2-13b-chat is a larger model fine-tuned for chat-based tasks. The results show that DONOD generalizes well across both cross-architecture and supervised fine-tuned models. As discussed in § 3.3.1, we select $l=32$, the lm_head layer (i.e., the output layer) of the Llama-3.1-8B-Instruct model, for our experiments. Additional details on training configurations can be found in Appendix  A.2. For ablations, we trained the Llama-3.1-8B-Instruct model with the SAT Math COT dataset pruned by the different configurations of alternative DONOD. The other training settings remain the same as the main settings.
4.2 Baselines
DONOD is compared with four baselines to evaluate its effectiveness. These baselines starts from random pruning, the full dataset. Additionally, to account for the model-intrinsic nature of DONOD, we compare it with two advanced methods that do not rely on external models: Instruction Following Difficulty (IFD) and Low-rank gradiEnt Similarity Search (LESS).
IFD, proposed by Li et al. (2024), selects data based on a novel metric, the Instruction Following Difficulty. This metric is measured by the ratio of losses for the same sample obtained with and without the instruction in the prompt. IFD assumes that samples where instructions significantly aid the model in producing the expected output should be prioritized. However, IFD has limitations. For datasets like LogiQA, where each sample is labeled with a single character, the oversimplified nature of the labels reduces IFD’s sensitivity and effectiveness. Furthermore, IFD relies on datasets in the Alpaca format, which requires a clear separation between instructions and input. For datasets like OpenAI’s, where the input consists of a single prompt, IFD is less efficient.
Regarding this, we also compare DONOD with LESS, a gradient-based method proposed by Xia et al. (2024). LESS selects influential data for instruction tuning by measuring gradient similarity between training examples and a target task. Unlike IFD, LESS has no specific dataset format requirements and does not overly rely on labels. However, LESS requires a clear target data distribution, meaning it necessitates a development set. This constraint limits its applicability in scenarios where such a set is unavailable.
Given the constraints of IFD and LESS, as well as the characteristics of the benchmarks (see Table  5), we train and evaluate IFD on the LogiQA-Train dataset and LESS on the SAT Math COT dataset. For a fair comparison, we follow the best settings reported in the respective papers and compare the results with the same amount or even less of data pruned by DONOD.
4.3 Main Results
Empirical Validation of Output Layer Focus
Figure 3: Rank of the top 5 layers by Frobenius norm delta after fine-tuning Llama-2-13B-Chat on SAT Math COT
Figure 4: Stability of DONOD across data proportions. Performance peaks at 20%-30% of data (optimal balance for Logic, Reading, and Overall Average), with Math scores steadily improving up to 30%. Notably, 10% DONOD-selected data outperforms full-dataset training (100%) in Logic and Overall Average, demonstrating robustness and efficiency.
To further justify our design choice in § 3.3.1, we analyze the sensitivity of weight changes across layers. Figure 4 ranks the Frobenius norm delta across layers after fine-tuning Llama-2-13B-Chat on the SAT Math COT dataset. The output layer (lm_head) shows the largest delta. We also tested on more models, details can be found in Appendix  A.8. By restricting DON/NOD computations to this layer, we retain of the total Frobenius signal (averaged across layers) while reducing computational costs to negligible. Comparing with complete backpropagation and compute $DON_{total}$ and ${NOD_{total}}$ across all layers for every single data sample, this only require backpropagation of the last layer of the model together with a constant time $DON$ and $NOD$ computation.
Stability of DONOD Across Data Proportions and Model Architecture
To evaluate the stability of DONOD under varying data regimes, we analyze its performance when trained on subsets of the dataset (10%–100%). As shown in Figure  4, DONOD demonstrates remarkable stability and efficiency, achieving competitive or superior performance even with significantly reduced data.
Notably, with only 10% of the data, DONOD surpasses the full-dataset (100%) baseline in Logic (37.43 vs. 37.12) and Overall Average (53.82 vs. 53.23), highlighting its robustness against data scarcity. At 20% data, the method reaches an optimal balance between data efficiency and model performance, achieving peak scores in Logic (38.96) and Reading (70.08), while maintaining strong Overall Average (56.01). This suggests DONOD effectively identifies high-quality samples that maximize learning signals with minimal noise. Similar results are obtained on Llama-3.2-3B-Instruct and Llama-2-13b-chat, detailed results can be found in Appendix  A.4.
A key finding is the steady improvement in mathematical reasoning as data increases to 30% (73.70), outperforming all other configurations. This aligns with DONOD’s ability to retain domain-critical samples that enhance specialized capabilities without overfitting. Beyond 30%, performance in most categories plateaus or declines slightly, indicating diminishing returns from additional data—a hallmark of DONOD’s precision in filtering redundant or harmful samples.
Table 1: Experimental results with domain-specific averages. The Non-Target column shows average performance excluding mathematical reasoning (target domain), revealing how methods generalize to other abilities. All values are percentages.
Configuration Logic (↑) Reading (↑) Math (↑) IFEval (↑) Non-Target (↑) All Avg (↑)
IFD 40% 39.31 68.26 64.18 63.03 56.87 55.04
Random 40% 38.32 69.72 61.17 59.89 55.98 53.15
ALL 100% 37.12 68.76 64.14 59.70 55.19 53.23
DONOD 20% 38.96 70.08 67.44 65.62 58.22 56.01
Table 2: Experimental results comparing LESS with other methods at 5% data selection. The Non-Target column shows average performance excluding logic reasoning (target domain), revealing how methods generalize to other abilities. All values are percentages.
LESS 5% 22.62 25.19 21.47 64.33 36.99 27.91
@@ 4.3 Main Results
Comparison with Baselines
As shown in Table 1 and  2, DONOD achieves superior overall performance across both experiments, despite using less data than baselines and requiring no development set. While minor trade-offs exist in some domains (e.g., logic at 5%: 20.91 vs. LESS’s 22.62), DONOD balances task-specific gains and cross-domain stability effectively.
Being more specific, with 20% data, DONOD surpasses IFD (40% data) by +0.97 All Avg (56.01 vs. 55.04) and outperforms the full dataset by +2.78. At 5% data, DONOD exceeds LESS (5%) by 0.95 All Avg (28.86 vs. 27.91) and improves over 100% data by +6.09, showcasing robustness to extreme sparsity. Also, our method shows a well target-generalization balance, in the target domain (math), DONOD achieves SOTA (67.44 at 20%, 24.22 at 5%), proving its ability to prioritize task-critical samples. Despite slight logic/reading trade-offs (e.g., 20.91 vs. LESS’s 22.62), DONOD retains best-in-class cross-domain performance (Non-Target: 58.22 at 20%, 39.15 at 5%), outperforming all baselines. Considering DONOD requires no development set and works seamlessly across both CoT (SAT Math) and non-CoT (LogiQA) datasets, such a minor trade-off should be acceptable.
4.4 Ablation Experiments Results
As shown in Table  4, We evaluate four configurations: (1) DON-only ranking, (2) NOD-only ranking, (3) DON + NOD combinations without TOPSIS (Weighted Sum and Pareto Front), and (4) the full method (DON + NOD + TOPSIS). Notably, we omit ablations combining TOPSIS with only one metric (e.g., DON + TOPSIS or NOD + TOPSIS) because TOPSIS inherently requires multiple criteria to resolve conflicting rankings. Applying TOPSIS to a single metric reduces it to a trivial ranking equivalent to using the metric alone, offering no performance gain. Results are presented in Table 3, revealing distinct roles for each component in balancing domain-specific performance and cross-domain generalization.
Configuration Logic (↑) Reading (↑) Math (↑) IFEval (↑) Non-Target (↑) All Avg (↑)
DON 36.19 68.12 58.15 66.17 56.83 52.38
@@ 4.4 Ablation Experiments Results
Using NOD alone improves target-domain performance (e.g., Math: 65.71% vs. DON’s 58.15%), as NOD prioritizes samples that can best boost model’s performance of the corresponding domain. However, this comes at the cost of reduced non-target generalization (55.13% vs. DON’s 56.83%), indicating over-specialization. Conversely, DON alone achieves better cross-domain robustness (Non-Target: 56.83%) by filtering samples with that harms the cross-domain generalization, though it lags in target-domain adaptation.
@@
Combining DON and NOD via naive strategies like Weighted Sum or Pareto Front fails to harmonize their strengths. Weighted Sum marginally improves Reading (69.15%) but degrades Math (63.74%) and Non-Target (53.67%) compared to single metrics. Pareto Front prioritizes stability (Non-Target: 57.37%) but sacrifices target-domain gains (Math: 57.82%). Both strategies highlight the challenge of balancing conflicting metrics without a principled integration method.
@@
Integrating DON, NOD, and TOPSIS achieves optimal balance. TOPSIS resolves metric conflicts by ranking samples based on their proximity to ideal DON/NOD values, improving target-domain accuracy (Math: 67.44%, +1.73% over NOD) while preserving cross-domain robustness (Non-Target: 58.22%, +1.39% over DON). The full method also dominates in overall performance (All Avg: 56.01%, +2.47% over Weight Sum), validating the necessity of TOPSIS for multi-objective optimization.
@@ 4.5 More Analytical Results
Computational Complexity
The algorithm’s computational complexity consists of two phases: (1) per-sample DON and NOD computation and (2) TOPSIS-based sample selection.
In the first phase, we perform a forward pass ($O(P)$ time per sample, where $P$ is the model’s parameter count), a backward pass restricted to the output layer ($O(O)$ time, $O$ being the output layer’s parameters), and compute Frobenius norms for weight updates ($O(O)$). Since $O\subset P$, the per-sample cost simplifies to $O(P)$, resulting in a total training complexity of $O(N\cdot P)$, where $N$ is the number of training samples.
Human Evaluation on Pruned Dataset
Figure 5: Illustration of the distribution of pruned dataset with DONOD and NODON (keep samples with the lowest TOPSIS scores, efficiently samples with high NOD and low DON).
To rigorously demonstrate the effectiveness of DONOD, a human evaluation was conducted. By reversing the selection process, where the top $N$ samples with the highest TOPSIS scores are retained, NODON generates a dataset containing the top samples that DONOD would filter out. As illustrated in Figure  5, these samples tend to cluster on the right side of the NOD-DON plane.
To further analyze the data pruned by our method, we examined samples from the NODON-pruned SAT Math COT datasets. The evaluation revealed several common categories of pruned samples: 1) Overthinking in response to easy questions, 2) Incomplete or partial answers, 3) Incorrect answers due to reasoning errors, 4) Incorrect answers resulting from errors in the question itself, 5) Overly complex or excessively difficult questions.
@@ 5 Conclusion
In this paper, we propose DONOD, a model-intrinsic dataset pruning framework for LLM SFT. Based on the model’s weight dynamics, our framework leverages dual complementary metrics, DON and NOD. DON captures the cumulative shifts in model weight magnitudes while NOD quantifies the instability caused by individual samples. Thus, DON prioritizes samples with stable cross-domain generalization , and NOD filters out noisy or unlearnable samples. Integrated by TOPSIS, our dataset pruning framework effectively balances the objective of maximizing generalization and minimizing harmful updates. Extensive experiment results demonstrate that our framework can significantly reduce training data volume (e.g., reducing 70%) with performance improvement compared to standard SFT. Meanwhile, the pruned datasets exhibit superior cross-architecture generalization. Datasets pruned by smaller models can generalize well on larger architectures, further enhancing the practicality.

### v2

@@ 4.1 Experiment Setup
Following (Ma et al. 2025), we construct our evaluation benchmark based on AGIEval (Zhong et al. 2024) and IFEval (Zhou et al. 2023). This benchmark is designed to be comprehensive and domain-orthogonal, assessing abilities in logical reasoning, mathematics, reading comprehension, and instruction following. By incorporating diverse datasets, the benchmark reflects real-world ad-hoc SFT scenarios, where the objective is to strengthen targeted model abilities rather than optimize for narrow or unrepresentative benchmarks.
[…]
LogiQA Train LESS 5% $22.62$ 25.19 $21.47$ $64.33$ $36.99$ $27.91$
@@
GSM8K ALL (100%) $33.43$ $70.74$ 47.99 $59.70$ $48.96$ 48.74
LESS 5% $34.74$ $71.85$ $25.78$ $60.63$ $50.46$ $44.97$
DONOD 30% 37.16 71.92 $33.69$ 71.16 52.74 $48.51$
Table 1: Experimental results with domain-specific averages on LogiQA Train and GSM8K. The Non-Target column shows average performance excluding logic reasoning or Math (target domain).
To comprehensively evaluate DONOD’s effectiveness across diverse domains, tasks, and data conditions, we select data that span a wide range of settings, including variations in domain (e.g., mathematics, logical reasoning, instruction following), task format (e.g., chain-of-thought, multiple-choice, fine-grained evaluation), and the presence or absence of validation sets. This design enables robust benchmarking under realistic and varied constraints. Specifically, we assess DONOD across the following settings:
  • SAT Math Chain-of-Thought (COT) (Davidson 2023) (math, COT, no validation set).
  • LogiQA-Train (Liu et al. 2020) (logical reasoning, multiple-choice, with validation set).
  • IFEval-like Data (Xu et al. 2024) (instruction-following, general, no validation set).
  • GSM8K (Cobbe et al. 2021) (math, COT, biased distribution vs. SAT Math and Aqua-RAT).
[…]
SAT Math COT IFD 40% $39.31$ $68.26$ $64.18$ $63.03$ $56.87$ $55.04$
ALL 100% $37.12$ $68.76$ $64.14$ $59.70$ $55.19$ $53.23$
Random 40 $38.32$ $69.72$ $61.17$ $59.89$ $55.98$ $53.15$
DONOD 30% 39.98 $67.62$ 73.70 $63.40$ $57.00$ 56.25
DONOD 20% $38.96$ 70.08 $67.44$ 65.62 58.22 $56.01$
IFEval-Like Data ALL (100%) $36.06$ $64.92$ $44.84$ $64.33$ $48.60$ $48.68$
IFD 40% $34.06$ 66.31 $34.04$ 71.90 $44.80$ $46.55$
DONOD 30% 36.14 $57.82$ 66.47 $46.77$ 53.47 49.01
Table 2: Experimental results with domain-specific averages on SAT Math COT and IFEval-Like Data. The Non-Target column shows average performance excluding mathematical reasoning or IFEval (target domain), revealing how methods generalize to other abilities. The All Avg column shows the average of all tasks in the benchmark. All values are percentages.
Models and Experiment Settings
We evaluate DONOD on a diverse set of instruction-tuned models to assess its generalizability across architectures and fine-tuning paradigms. Specifically, we consider: (1) LLaMA-3.2-3B-Instruct (Meta AI 2024), a lightweight model optimized for instruction-following tasks; (2) LLaMA-3.1-8B-Instruct (Grattafiori et al. 2024), a mid-sized model widely used in recent instruction-tuning studies; (3) LLaMA-2-13B-Chat (Touvron et al. 2023), a larger model trained with conversational objectives; and (4) Qwen 2.5-7B-Instruct (Team 2024), a model from a distinct architecture family, differing in tokenizer, training data, and parameterization. This ensures a comprehensive evaluation of DONOD under varied model designs and training strategies, i.e., cross-architecture generalization and supervised fine-tuning performance. As described in Section 3.2, we focus on the output layer of the Llama-3.1-8B-Instruct model for our experiment.
4.2 Main Results
ALL (100%) 43.62 73.99 74.57 56.01 57.87 58.90
DONOD 20% 41.95 68.28 79.86 58.23 56.15 59.09
DONOD 30% 42.57 71.15 77.07 61.74 58.48 59.33
Table 3: Cross-architecture generalization of selecting data with Llama-3.1-8B-Instruct and fine-tuning on Qwen-2.5-7B-Instruct.
Comparison with State-of-the-arts
As shown in Table 1 and Table 2, DONOD consistently outperforms others while using significantly less data across nearly all benchmarks, e.g., 20-30%, and requiring no development set. It achieves state-of-the-art performance in core reasoning tasks such as math and logic, outperforming full-data baselines in both target-domain accuracy and cross-domain generalization.
Compared to IFD and LESS, DONOD demonstrates better balance between accuracy and efficiency, achieving higher overall and non-target averages under constrained data budgets. Notably, it exhibits strong cross-task and cross-domain transferability, without relying on task-specific tuning or heuristics, and remains competitive even in challenging settings, such as reading comprehension.
These results highlight DONOD’s ability to retain or exceed full-data performance while substantially reducing training cost, which offers a robust and scalable solution for data-efficient LLM fine-tuning.
ALL (100%) 30.57 54.68 23.38 29.94 38.39 35.28
DONOD 20% 31.52 55.58 26.89 29.39 38.83 36.57
DONOD 30% 31.30 54.61 23.39 29.57 38.49 35.53
Table 4: Cross-architecture generalization of selecting data with Llama-3.1-8B-Instruct and fine-tuning on Llama-2-13b-chat.
[…]
Main Results DON $36.19$ $68.12$ $58.15$ 66.17 $56.83$ $52.38$
@@ 4.2 Main Results
4.3 Cross-Architecture Generalization
To evaluate cross-architecture generalization, we select data using Llama-3.1-8B-Instruct and fine-tune on both Llama-2-13b-chat and a structurally different model, Qwen-2.5-7B-Instruct. As shown in Table 3 and Table 4, DONOD-selected subsets (20% and 30%) not only retain but surpass the full-data baseline in overall performance. These results suggest that despite differences in model size and structure, language models exhibit consistent rankings in perceiving instruction difficulty. In (Li et al. 2024a), this consistency is demonstrated through metrics like perplexity and Instruction-Following Difficulty scores, which show strong rank correlations across models of different sizes. As a result, smaller models like GPT-2 can effectively filter instruction data for much larger models, such as LLaMA2-7B or GPT-4.
DONOD leverages this same principle by selecting data based on intrinsic parameter-level signals. These signals identify samples that are universally useful across architectures. In these experiments, we observe that pruning data using DONOD on a smaller model (e.g., Llama-3.1-8B) leads to better performance when fine-tuning larger or different models (e.g., Llama-2-13B and Qwen-2.5-7B), echoing the weak-to-strong transfer effect. These results demonstrate strong transferability of selected samples, highlighting DONOD’s practical utility in real-world settings.
4.4 Robustness in Identifying Noise
To evaluate DONOD’s robustness in identifying noisy data, we conduct a controlled experiment on the SAT Math CoT dataset using the LLaMA 3.1 model. We start with the top 20% of samples originally selected by DONOD.
To simulate real-world data imperfections, we introduce controlled noise by randomly masking words in the labels of these clean samples, which mimics subtle corruption or annotation errors. These perturbed samples are then reintegrated into the full dataset, creating a new training pool containing embedded noisy instances. We reapply DONOD to this dataset to select a new top 20% subset based on the updated DON and NOD values. We assess sensitivity to noise by measuring the overlap between the original and newly selected top 20%. The result shows a drop to only 38.7% overlap, indicating that DONOD successfully identifies and filters out many of the newly corrupted samples.
This experiment highlights DONOD’s strong responsiveness to fine-grained label corruption and its ability to dynamically adapt selection criteria. Such robustness is critical in real-world scenarios where large-scale instruction data, often web-scraped or weakly supervised, is prone to noise and inconsistency.
4.5 Ablation Study
We conduct experiments using LLaMA-3.1-8B-Instruct on the SAT Math CoT dataset to analyze the effect of different components in DONOD. As shown in Table  5, we evaluate four configurations: (1) DON-only ranking, (2) NOD-only ranking, (3) DON + NOD combinations without TOPSIS (Weighted Sum and Pareto Front), and (4) the full method (DON + NOD + TOPSIS). Notably, we omit ablations combining TOPSIS with only one metric (e.g., DON + TOPSIS) because TOPSIS inherently requires multiple criteria to resolve conflicting rankings.
@@ 4.5 Ablation Study
When using NOD alone, we observe stronger performance in the target domain (e.g., Math: 65.71%) compared to DON (58.15%), as NOD emphasizes samples that lead to significant, localized parameter updates, favoring task-specific learning. However, this comes at the cost of reduced generalization to non-target domains, suggesting overfitting to task-specific patterns.
In contrast, DON alone yields better cross-domain generalization by favoring samples that contribute to stable and smooth parameter updates, which better preserve generalizable knowledge. However, this approach underperforms in task-specific adaptation, reflecting a trade-off between stability and specificity.
@@
Simply combining DON and NOD via weighted sum or Pareto Front fails to effectively reconcile the competing goals of DON and NOD. The weighted sum approach marginally improves Reading but degrades Math and Non-Target compared to using DON or NOD individually. Similarly, the Pareto Front strategy emphasizes cross-domain stability but sacrifices gains in the target domain. These results underscore the difficulty of balancing conflicting criteria without a principled integration mechanism.
@@
The proposed DONOD achieves the best trade-off between domain-specific performance and cross-domain generalization. By ranking samples based on their proximity to the ideal combination of DON and NOD, TOPSIS effectively resolves conflicts between the two objectives. This leads to improvements in both target-domain accuracy and cross-domain robustness. The overall average performance also improves significantly, validating the effectiveness and importance of using TOPSIS as a principled strategy for multi-objective optimization.
@@ 4.6 More Analytical Results
Figure 3: Rank of the top 5 layers by Frobenius norm delta after fine-tuning Llama-2-13B-Chat on SAT Math COT.
Figure 4: Stability of DONOD across data proportions.
Validation of Output Layer Focus
To further justify our design choice in Section 3.2, we analyze the sensitivity of weight changes across layers. Ranking the Frobenius norm delta across layers after fine-tuning Llama-2-13B-Chat on the SAT Math COT dataset, as shown in Figure 3, the output layer exhibits the largest delta, indicating its heightened responsiveness to task-specific supervision. By restricting DON and NOD to this layer, we retain a representative signal of overall weight dynamics (closely aligned with the layer-wise average Frobenius norm) while significantly reducing computational overhead. In contrast to full backpropagation and computing $DON_{total}$ and $NOD_{total}$} across all layers for every data sample, our approach only requires backpropagation through the final layer, along with a constant time $DON$ and $NOD$ computation, yielding a highly efficient yet effective approximation.
Stability of DONOD Across Data Proportions
To evaluate the stability of DONOD under varying data regimes, we analyze the performance when trained on the selected datasets (10%–100%). As shown in Figure 4, DONOD demonstrates remarkable stability and efficiency, achieving competitive or better performance even with significantly reduced data.
Notably, with only 10% of the data, DONOD surpasses the full-dataset baseline in Logic and Overall Average, highlighting its robustness against data scarcity. At a selection ratio of 20%, our method achieves an optimal balance between data efficiency and model performance, achieving peak scores in Logic and Reading, while maintaining a strong overall average. This suggests that DONOD effectively identifies high-quality samples that maximize learning signals with minimal noise.
Human Evaluation on Selected Datasets
Figure 5: Illustration of the distribution of the selected dataset with DONOD and NODON.
To rigorously demonstrate the effectiveness of DONOD, a human evaluation is conducted. By reversing the selection process, where the top $N$ samples with the highest TOPSIS scores are retained, NODON generates a dataset containing the top samples that DONOD would filter out. As illustrated in Figure 5, these samples tend to cluster on the right side of the NOD-DON plane.
To further analyze the data pruned by our method, we then examine the samples from the NODON-pruned SAT Math COT datasets. The evaluation reveals several common categories of pruned samples: 1) Overthinking in response to easy questions, 2) Incomplete or partial answers, 3) Incorrect answers due to reasoning errors, 4) Incorrect answers resulting from errors in the question itself, and 5) Overly complex or excessively difficult questions. These findings suggest that DONOD effectively filters out unhelpful or misleading samples, thereby improving instruction data quality and enhancing the overall robustness and generalization of fine-tuned LLMs.
@@ 5 Conclusion
In this paper, we propose DONOD, a model-intrinsic dataset pruning framework to enhance LLM fine-tuning efficiency without sacrificing model performance. By leveraging weight dynamics, our method selects high-quality data and suppresses the selection of noisy or uninformative data via dual complementary metrics, DON and NOD. These metrics are integrated via TOPSIS, enabling a principled trade-off between maximizing generalization and minimizing harmful updates. Experiments show that DONOD can reduce training data volume by up to 70% while outperforming standard supervised fine-tuning, achieving superior training acceleration. Notably, datasets selected by smaller models also generalize well when used to fine-tune other LLM architectures, underscoring the framework’s scalability and practicality for real-world LLM pipelines. We hope DONOD inspires further research on data selection for LLM training from a model-intrinsic perspective and believe our method will serve as a promising dataset optimization tool for the community, enabling enhanced data-centric LLM training pipelines.